Find the sum of two consecutive positive even integers such that the square of the smaller is 4 more than 10 times the larger.
I need to know the procces so show steps please.

Answer by mananth(16946)

Find the sum of two consecutive positive even integers such that the square of the smaller is 4 more than 10 times the larger.
..
let the numbers be x & x+2
the square of the smaller = x^2
10 times larger = 10(x+2)
4 more
10(x+2)+4
x^2=10(x+2)+4
x^2=10x+20+4
x^2=10x+24
x^2-10x-24=0
x^2-12x+2x-24=0
x(x-12)+2(x-12)=0
(x-12)(x+2)
so the number is x=12 which is positive
the other number =12+2= 14
12+14=26 their sum
...
m.ananth@hotmail.ca